# Neuromorphic Overparameterisation: Generalisation and Few-Shot Learning in Multilayer Physical Neural Networks

### v4

@@ Abstract
The networking approach is an effective method of improving readout dimensionality. Combining distinct systems is beneficial as the breadth of dynamics is likely to be higher, but networking the same physical system can produce overparameterisation. Enriched dynamics and increased effective readout dimensionality can be achieved by operating the physical system in different dynamic regimes, for example, by changing external conditions to access different dynamics[45]. Alternatively, by varying the input sequence, either through a random mask (a common technique known as ‘virtual nodes’) or by feeding the output of one node as input to another, a diverse parallel / series network can be created. Even in the limit of a single physical system with one readout per time-step, overparameterisation can be achieved through a combination of parallel and series networking, provided that the internal dynamics are rich enough.
@@
Figure 4 a) shows the system prediction when predicting (left-to-right) the , and NARMA-processed Mackey-Glass signals. The PNN is trained on just the first 50 data points of the signal, highlighted by black circles. The PNN is able to learn the underlying system dynamics and provide good predictions demonstrating the power and adaptability of the overparameterised regime. We note that the MSE’s achieved are are comparable to those in Figure 2 e,g) when using feature selection with the PNN. As such, this meta-learning can achieve strong results with a 75 reduction in training set size (50 training data points here vs 200 previously used).
To further showcase the computational capabilities of the PNN, we now demonstrate few-shot learning where the ‘seen’ training data are sparsely distributed throughout the target dataset, representing for instance a very low sampling rate of a physical input sensor in an edge-computing use-case. In Figure 4 b), the system is driven by a sinusoidal input and asked to predict a target of the form i.e. simultaneously predict amplitude, frequency and phase changes - a task that requires a range of temporal dynamics in the network, often used as a meta-learning benchmark task[31]. The values and are sampled randomly at the beginning of each task from continuous uniform distributions (details in the methods section Task selection). The goal is to train the system to generalise to a set of and values, and then rapidly adapt to a new task with a limited set of sparse training points.
@@
Figure 4 b) and c) show predictions of three example tasks for the overall target and two example sub-components from each task respectively. The network sees just 15 data points (highlighted by grey circles in panel b)) throughout the entire process. In Figure 4 b), the target (dashed black line) and predicted values after updating the generalised matrix a number of times are shown. At update 0, the predicted response does not match the target waveform as expected. As the network updates, the error between the prediction and target reduce. Further example tasks are provided in supplementary note 12.
Despite the limited information and the high variability of tasks, the network learns the underlying sinusoidal components and can adapt to different targets with the partial information available. To support this claim of generalisation, Figure 4 d) and e) show the average error and the average variance of the error respectively, calculated over 500 different tasks for various few-shot train lengths and number of updates. The error decreases as the number of updates and available data points is increased as expected, with strong performance observed for as little as 10 training data points. Crucially, the variance of the error across all tasks is low, demonstrating strong generalisation. Finally, Figure 4 f) reports an example of prediction obtained when meta-learning is applied on a single reservoir, which fails completely at the task.
@@ Discussion
We now propose and benchmark potential ways to translate nanomagnetic arrays to device level (extended analysis and discussion is provided in Supplementary note 13). We assume a nanomagnetic array with dimensions of 5 5 (100 elements). By patterning the array on top of a current-carrying microstrip, array-specific Oersted fields can be generated. For a 5 5 array patterned on top of an Cu microstrip, with width = 5 and thickness = 50 nm in height. The current required to produce a 25 mT Oersted field is 200 mA [46] (see Supplementary note 13 for calculation details). For Cu resistivity of 16.8 [47], the resistance is R = 0.336 consuming a minimum power P = 13.4 mW. For a pulse time of 1 ns, this equates to an energy of 13.4 pJ per input. Nanomagnet coercivity and therefore power consumption can be reduced by fabricating thinner nanomagnets.
@@
Instead, we can fabricate arrays in contact with a high spin-Hall angle material to switch macrospins via spin-orbit torque[48]. This can be achieved with Ta thickness and magnetic layer thicknesses of around 5 nm and 1-2 nm respectively[48] and current densities on the order of 1011 . Whilst reducing thickness reduces dipolar coupling, arrays of nanomagnets remain highly correlated at these thicknesses[49, 50] and will retain the collective processing that thicker nanomagnets exhibit. For a 5 5 Ta strip with 5 nm thickness, the corresponds to a current of 2.5 mA. For Ta resistivity of 131 [51], RTa = 26.9 and P = 0.17 mW. Here, switching times can be as low as 250 ps [52] giving input energies of 42.5 fJ. Spin-orbit torques decrease linearly with increasing ferromagnet thickness [53]. For 20 nm thick nanomagnets used in this work, the current density required to switch an element will be on the order of 10 - 40 (based on micromagnetic simulations) due to thicker elements and higher coercive fields, giving currents in the range of 25 - 100 mA and powers in the range of 16.8 - 269 mW. Limited spin-diffusion in ferromagnets may further increase or even prevent switching in thicker elements. As such, for thicker nanomagnets, Oersted field switching may be preferable.
@@
Microwave readout can be implemented either sequentially (frequency-swept RF source) or in parallel (multiple channels simultaneously) depending on the technique. For sequential readout out (as implemented in this work), we can utilise spin-torque FMR which converts magnetisation resonance to DC voltage. Assuming microwave powers in the range of 1 - 10 mW, the generated voltage at resonance is on the order 1 - 10 which can be amplified with a CMOS low noise high gain amplifier. Readout time per channel is theoretically limited to the time taken to reach steady precession ( 14 ns[54]). As the generated voltage is DC, spectral information is lost, hence each channel must be recorded sequentially, giving a measurement time of 2.6 s and energy of 2.8 nJ (supplementary note 13), dominating over the input speeds. Alternatively, we can pass mixed frequency rf-signals through a device patterned on a co-planar waveguide and detect the microwave absorption, for example via an rf noise source. Here, the readout power is higher (150 mW) but speed and energy are reduced (14 ns and 2.1 nJ respectively, supplementary note 13). The output signal can, for example, be sent to multiple spin-torque oscillators tuned to different frequencies serving as rf-diodes[55, 54] allowing parallel detection of multiple channels at ns timescales. Weight multiplication can be achieved by passing output signals to memristor cross-bar arrays which routinely operate with powers[56]. As such, the power consumption, energy and time required to perform an operation of our proposed device is on the order of 1 - 160 mW, 1.5 - 2800 ns and 2.1 - 2.9 nJ depending on the input and readout method used. At the lower end, this is within the power range for ‘battery-less’ operation using energy harvesting technology[57, 58]. For PNNs, nodes are measured sequentially hence the total operating power remains unchanged, and time and energy scale linearly with the number of nodes. Simultaneous measurement of physical nodes would increase power and reduce time. Supplementary note 13 provides further comparison to CMOS hardware by calculating the number of FLOPs to update an echo-state network (the closest software analogue to nanomagnetic hardware). Our calculations indicate that, for large echo-state networks, the proposed devices have potential to operate at lower energy costs that conventional hardware.
Conclusion
@@
We mitigate the limitations of physical reservoir computing by engineering networks of physical reservoirs with distinct properties. We have engineered multiple nanomagnetic reservoirs with varying internal dynamics, evaluating their computational metrics and performance across a broad benchmark taskset. Our results highlight the computational performance gained from enriched state spaces, applicable across a broad range of neuromorphic systems. PW outperforms MS by up to 31.6, demonstrating careful design of system geometry and dynamics is critical, with computational benefits available via physical system design optimisation.

### v5

@@ Results
The networking approach is an effective method of improving readout dimensionality. Combining distinct systems is beneficial as the breadth of dynamics is likely to be higher, but networking the same physical system can produce overparameterisation. Enriched dynamics and increased effective readout dimensionality can be achieved by operating the physical system in different dynamic regimes, for example, by changing external conditions to access different dynamics[46]. Alternatively, by varying the input sequence, either through a random mask (a common technique known as virtual nodes) or by feeding the output of one node as input to another, a diverse parallel / series network can be created. Even in the limit of a single physical system with one readout per time-step, overparameterisation can be achieved through a combination of parallel and series networking, provided that the internal dynamics are rich enough.
@@
Figure 4 a) shows the system prediction when predicting (left-to-right) the , and NARMA-processed Mackey-Glass signals. The PNN is trained on just the first 50 data points of the signal, highlighted by black circles. The PNN is able to learn the underlying system dynamics and provide good predictions demonstrating the power and adaptability of the overparameterised regime. We note that the MSE’s achieved are are comparable to those in Figures 2 e,g) when using feature selection with the PNN. As such, this meta-learning can achieve strong results with a 75 reduction in training set size (50 training data points here vs 200 previously used).
To further showcase the computational capabilities of the PNN, we now demonstrate few-shot learning where the seen training data are sparsely distributed throughout the target dataset, representing for instance a very low sampling rate of a physical input sensor in an edge-computing use-case. In Figure 4 b), the system is driven by a sinusoidal input and asked to predict a target of the form ỹ(t)= ( + ) i.e. simultaneously predict amplitude, frequency and phase changes - a task that requires a range of temporal dynamics in the network, often used as a meta-learning benchmark task[31]. The values and are sampled randomly at the beginning of each task from continuous uniform distributions (details in the Methods section, Task Selection). The goal is to train the system to generalise to a set of and values, and then rapidly adapt to a new task with a limited set of sparse training points.
@@
Figures 4 b) and c) show predictions of three example tasks for the overall target and two example sub-components from each task respectively. The network sees just 15 data points (highlighted by grey circles in panel b)) throughout the entire process. In Figure 4 b), the target (dashed black line) and predicted values after updating the generalised matrix a number of times are shown. At update 0, the predicted response does not match the target waveform as expected. As the network updates, the error between the prediction and target reduce. Further example tasks are provided in Supplementary Note 12.
Despite the limited information and the high variability of tasks, the network learns the underlying sinusoidal components and can adapt to different targets with the partial information available. To support this claim of generalisation, Figures 4 d) and e) show the average error < MSE >, and variance of the error, < 2 MSE >, respectively, calculated over 500 different tasks for various few-shot train lengths and number of updates. The error decreases as the number of updates and available data points is increased as expected, with strong performance observed for as little as 10 training data points. Crucially, the variance of the error across all tasks is low, demonstrating strong generalisation. Finally, Figure 4 f) reports an example of prediction obtained when meta-learning is applied on a single reservoir, which fails completely at the task.
@@ Discussion
We now propose and benchmark potential ways to translate nanomagnetic arrays to device level (extended analysis and discussion is provided in Supplementary Note 13). We assume a nanomagnetic array with dimensions of m (100 elements). By patterning the array on top of a current-carrying microstrip, array-specific Oersted fields can be generated. For a array patterned on top of an Cu microstrip, with width = and thickness = 50 nm in height. The current required to produce a 25 mT Oersted field is 200 mA [47] (see Supplementary Note 13 for calculation details). For Cu resistivity of 16.8 n m[48], the resistance is R = 0.336 consuming a minimum power P = 13.4 mW. For a pulse time of 1 ns, this equates to an energy of 13.4 pJ per input. Nanomagnet coercivity and therefore power consumption can be reduced by fabricating thinner nanomagnets.
@@
Instead, we can fabricate arrays in contact with a high spin-Hall angle material to switch macrospins via spin-orbit torque[49]. This can be achieved with Ta thickness and magnetic layer thicknesses of around 5 nm and 1-2 nm respectively[49] and current densities on the order of 1011 Am-2. Whilst reducing thickness reduces dipolar coupling, arrays of nanomagnets remain highly correlated at these thicknesses[50, 51] and will retain the collective processing that thicker nanomagnets exhibit. For a Ta strip with 5 nm thickness, the corresponds to a current of 2.5 mA. For Ta resistivity of 131 nm[52], RTa = 26.2 and P = 0.164 mW. Here, switching times can be as low as 250 ps [53] giving input energies of 41 fJ. Spin-orbit torques decrease linearly with increasing ferromagnet thickness [54]. For 20 nm thick nanomagnets used in this work, the current density required to switch an element will be on the order of 10 - 40 (based on micromagnetic simulations) due to thicker elements and higher coercive fields, giving currents in the range of 25 - 100 mA and powers in the range of 16.8 - 269 mW. Limited spin-diffusion in ferromagnets may further increase or even prevent switching in thicker elements. As such, for thicker nanomagnets, Oersted field switching may be preferable.
@@
Microwave readout can be implemented either sequentially (frequency-swept RF source) or in parallel (multiple channels simultaneously) depending on the technique. For sequential readout out (as implemented in this work), we can utilise spin-torque FMR which converts magnetisation resonance to DC voltage. Assuming microwave powers in the range of 1 - 10 mW, the generated voltage at resonance is on the order 1 - 10 µV which can be amplified with a CMOS low noise high gain amplifier. Readout time per channel is theoretically limited to the time taken to reach steady precession ( 14 ns[55]). As the generated voltage is DC, spectral information is lost, hence each channel must be recorded sequentially, giving a measurement time of 2.6 µs and energy of 2.8 nJ (Supplementary Note 13), dominating over the input speeds. Alternatively, we can pass mixed frequency RF-signals through a device patterned on a co-planar waveguide and detect the microwave absorption, for example via an RF noise source. Here, the readout power is higher (150 mW) but speed and energy are reduced (14 ns and 2.1 nJ respectively, Supplementary Note 13). The output signal can, for example, be sent to multiple spin-torque oscillators tuned to different frequencies serving as RF-diodes[56, 55] allowing parallel detection of multiple channels at ns timescales. Weight multiplication can be achieved by passing output signals to memristor cross-bar arrays which routinely operate with nW - µW powers[57]. As such, the power consumption, energy and time required to perform an operation of our proposed device is on the order of 1 - 160 mW, 1.5 - 2,800 ns and 2.1 - 2.9 nJ depending on the input and readout method used. At the lower end, this is within the power range for battery-less operation using energy harvesting technology[58, 59]. For PNNs, nodes are measured sequentially hence the total operating power remains unchanged, and time and energy scale linearly with the number of nodes. Simultaneous measurement of physical nodes would increase power and reduce time. Supplementary Note 13 provides further comparison to CMOS hardware by calculating the number of FLOPs to update an echo-state network (the closest software analogue to nanomagnetic hardware). Our calculations indicate that, for large echo-state networks, the proposed devices have potential to operate at lower energy costs that conventional hardware.
@@
To conclude, here we mitigate the limitations of physical reservoir computing by engineering networks of physical reservoirs with distinct properties. We have engineered multiple nanomagnetic reservoirs with varying internal dynamics, evaluating their computational metrics and performance across a broad benchmark taskset. Our results highlight the computational performance gained from enriched state spaces, applicable across a broad range of neuromorphic systems. PW outperforms MS by up to 31.6, demonstrating careful design of system geometry and dynamics is critical, with computational benefits available via physical system design optimisation.
@@
Methods
The Methods section is organised as follows: Experimental Methods includes the fabrication of samples, measurement of FMR response and the details of implementing reservoir computing and interconnecting arrays. Following this we discuss the tasks chosen and how they are evaluated in Task selection. Learning algorithms then provides a detailed description of the learning algorithms used in this work.
Experimental Methods
Nanofabrication
Artificial spin reservoirs are fabricated via electron-beam lithography liftoff method on a Raith eLine system with PMMA resist. 25 nm Ni81Fe19 (permalloy) is thermally evaporated and capped with 5 nm Al2O3. For WM, a staircase subset of bars are increased in width to reduce its coercive field relative to the thin subset, allowing independent subset reversal via global field. For PW, a variation in widths are fabricated across the sample by varying the electron beam lithography dose. Within a write-field, the bar dimensions remain constant. The flip-chip FMR measurements require mm-scale nanostructure arrays. Each sample has dimensions of roughly 3 2 mm2. As such, the distribution of nanofabrication imperfections termed quenched disorder is of greater magnitude here than typically observed in studies on smaller artificial spin systems, typically employing 10-100 micron-scale arrays. The chief consequence of this is that the Gaussian spread of coercive fields is over a few mT for each bar subset. Smaller artificial spin reservoir arrays have narrower coercive field distributions, with the only consequence being that optimal applied field ranges for reservoir computation input will be scaled across a corresponding narrower field range, not an issue for typical 0.1 mT or better field resolution of modern magnet systems.
Magnetic Force Microscopy Measurement
Magnetic force micrographs are produced on a Dimension 3100 using commercially available normal-moment MFM tips.
Ferromagnetic Resonance Measurement
Ferromagnetic resonance spectra are measured using a NanOsc Instruments cryoFMR in a Quantum Design Physical Properties Measurement System. Broadband FMR measurements are carried out on large area samples mounted flip-chip style on a coplanar waveguide. The waveguide is connected to a microwave generator, coupling RF magnetic fields to the sample. The output from waveguide is rectified using an RF-diode detector. Measurements are done in fixed in-plane field while the RF frequency is swept in 20 MHz steps. The DC field are then modulated at 490 Hz with a 0.48 mT RMS field and the diode voltage response measured via lock-in. The experimental spectra show the derivative output of the microwave signal as a function of field and frequency. The normalised differential spectra are displayed as false-colour images with symmetric log colour scale.
Data Input and Readout
Reservoir computing schemes consist of three layers: an input layer, a hidden reservoir layer, and an output layer corresponding to globally-applied fields, the nanomagnetic reservoir and the FMR response respectively. For all tasks, the inputs are linearly mapped to a field range spanning 35 - 42 mT for MS, 18 - 23.5 mT for WM and 30-50 mT for PW, with the mapped field value corresponding to the maximum field of a minor loop applied to the system. In other words, for a single data point, we apply a field at + then -. After each minor loop, the FMR response is measured at the applied field - between 8 - 12.5 GHz, 5 - 10.5 GHz and 5 - 10.5 GHz in 20 MHz steps for MS, WM and PW respectively. The FMR output is smoothed in frequency by applying a low-pass filter to reduce noise. Eliminating noise improves computational performance[5]. For each input data-point of the external signal , we measure distinct frequency channels and take each channel as an output. This process is repeated for the entire dataset with training and prediction performed offline.
Interconnecting Arrays
When interconnecting arrays, we first input the original Mackey-Glass or sinusoidal input into the first array via the input and readout method previously described. We then analyse the memory and nonlinearity of each individual frequency output channel (described later). A particular frequency channel of interest is converted to an appropriate field range. The resulting field sequence then applied to the next array via the computing scheme previously described. This process is then repeated for the next array in the network. The outputs from every network layer are concatenated for learning.
Task Selection
Throughout this manuscript, we focus on temporally-driven regression tasks that require memory and nonlinearity. Considering a sequence of T inputs = , the physical system response is a series of observations = across time. These observations can be gathered from a single reservoir configuration as in Figure 1, or can be a collection of activities from multiple reservoirs, in parallel or inter-connected as in Figure 2. In other words, the response of the system at time t is the concatenation of the outputs of the different reservoirs used in the architecture considered. The tasks faced can be divided into five categories:
Sine Transformation Tasks. The system is driven by a sinusoidal periodic input and asked to predict different transformations, such as . The inputs are chosen to have 30 data points per period of the sinusoidal wave, thus with . The total dataset size is 250 data points. If the target is symmetric with respect to the input, the task only requires nonlinearity. If the target is asymmetric, then both nonlinearity and memory are required.
Mackey-Glass Forecasting. The Mackey-Glass time-delay differential equation takes the form and is evaluated numerically with = 0.2, n = 10 and = 17. Given as external varying input, the desired outputs are , corresponding to the future of the driving signal at different times. We use 22 data points per period of the external signal for a total of 250 data points. This task predominantly requires memory as constructing future steps requires knowledge of the previous behaviour of the input signal.
Non-Linear Auto-Regressive Moving Average Tasks. Non-linear auto-regressive moving average (NARMA) is a typical benchmark used by the reservoir computing community. The definition of the x-th desired output is , where the constants are set to A = 0.3, B = 0.01, C = 2, D = 0.1. The input signal is the Mackey-Glass signal, where the variable is introduced to account for a possible temporal shift of the input. For , is the application of NARMA on the Mackey-Glass signal at the current time t, while for , is the result of NARMA on the input signal delayed by ten time steps in the future. The index can instead vary between one, defining a task with a single temporal dependency, and fifteen, for a problem that requires memory of fifteen inputs. Varying and , we can define a rich variety of tasks with different computational complexity.
Evaluations of Memory-Capacity and Nonlinearity. Memory capacity and nonlinearity are metrics frequently used for the characterization of the properties of a physical device. While these metrics do not constitute tasks in the common terminology, we include them in this section for simplicity of explanation. Indeed, we use the same training methodology to measure MC and NL as in the other tasks faced. We evaluate these metrics with the Mackey-Glass time-delay differential equation as input. This gives results that are correlated to conventional memory-capacity and nonlinearity scores, with some small convolution of the input signal - negligible for our purposes of relatively assessing artificial spin reservoirs and designing network interconnections.
For memory-capacity the desired outputs are , corresponding to the previous inputs of the driving signal at different times. To avoid effects from the periodicity of the input signal, we set =8. The R2 value of the predicted and target values is evaluated for each value of where a high R2 value means a good linear fit and high memory and a low R2 value means a poor fit a low memory. The final memory-capacity value is the sum of R2 from = 0 to = 8.
Memory capacity and nonlinearity can be calculated using a single multiple frequency channels. For the single channel analysis, we perform the same calculations, but using just a single FMR channel.
Frequency Decomposition, a Few-Shot Learning Task. The network is driven by a sinusoidal input and needs to reconstruct a decomposition of a temporal varying signal in the form of = , where . The values of and are randomly sampled at the beginning of each task from uniform distributions. In particular, and . The output layer is composed of nodes, and the system is asked to predict a target after observing the values of over K data points i.e. time steps. The value of K adopted for the examples of Figures 4 b) and c) is fifteen, but the network reports good performance even with (Figures 4 d) and e). To face this challenging task, we use the PNN in the overparameterised regime and a meta-learning algorithm to quickly adapt the read-out connectivity. Details of the meta-learning approach are given below in the Meta-learning section of the Methods.
Learning Algorithms
The type of learning algorithm we use to select features and train the networks varies throughout the manuscript.
For comparisons between different systems (Figures 1 2), training is accomplished through a features selection algorithm (discussed below) and optimisation of the read-out weights .
When exploring the effects of overparameterisation, we randomly select features and then optimise for those features.
We call the representation at which the read-out weights operate, and we define the output of the system as . The vector simply contains a subset of features of defined via the feature selection algorithm or randomly picked as in the dimensional study on overparameterisation.
In this setting, optimisation of is achieved with linear or ridge regression, which minimises the error function .
Feature Selection
The dimensionality of an observation can vary depending on the architecture considered, spanning from dimensions when using a single arrays to 14,000 for the PNN of Figure 2. The high readout dimensionality allows better separability of input data, however, high dimensional spaces constitute a challenge due to overfitting issues. As such, learning over a high-dimensional features’ space with few data points constitutes a challenge and opportunity for physically defined reservoirs. For this reason, we design a feature-selection methodology to avoid overfitting and to exploit the computational abilities of architectures with varying complexity (Figure 5, Algorithms 1,2). The methodology adopted can be at first described as a cross-validation (inner validation loop) of a cross-validation approach (outer validation loop), where the outer cross-validation is used to accurately evaluate the performance and the inner loop is used to perform feature-selection (Figure 6). For each split, feature selection is accomplished by discarding highly correlated features and through an evolutionary algorithm. The independent parts of this methodology are known, but the overall procedure is unique and can give accurate performance measurements for our situation, where we have a suite of systems with varying dimensionality to compare over limited data.
Considering a specific split of the outer validation loop (Figure 6), where we select a validation set and a test set (comprising of the of the data each, i.e. 25 data points), we perform cross-validations on the remaining data to optimise hyperparameter values through grid-search. In this inner validation loop, each split corresponds to a test set (comprising again of the of the remaining data, without and ), where changing means to select a different test-split in the inner loop based on the -th original split of the outer validation (Figure 6). The remaining data, highlighted in green in Figure 6, are used for training to optimise the read-out weights and minimise the error function through ridge-regression previously described.
At this stage, we perform a grid-search methodology on hyperparameters and which control directly and indirectly the number of features being adopted for training (Algorithm 1). The hyperparameter acts as a threshold on the correlation matrix of the features. Simply, if the correlation among two features exceeds the specific value of considered, one of these two features is removed for training (and testing). The idea behind this method is to discard features that are highly correlated, since they would contribute in a similar way during training. This emphasises diversity in the reservoir response. The hyperparameter is the penalty term in ridge-regression. Higher values of lead to a stronger penalisation on the magnitude of the read-out weights. As such, can help prevent overfitting and controls indirectly the number of features being adopted. We should use a high value of if the model is more prone to overfitting the training dataset, a case that occurs when the number of features adopted is high. Calling the error computed on the test set with weights optimised on the corresponding training data ( in Algorithm 1 and with hyperparameter values and respectively, we select the values of the hyperparameters that correspond to the minimum average error over the test sets in the inner validation loop. Otherwise stated, we select the optimal and for the -th split in the outer loop from the test average error in the inner cross-validation as . This methodology permits to find and that are not strongly dependent on the split considered, while maintaining the parts of the dataset and unused during training and hyperparameter selection. The sets and correspond to the values explored in the grid-search. In our case, and . Repeating this procedure for each split of the outer loop, we found the optimal and , for . This concludes the hyperparameter selection algorithm described in Algorithm 1. Selection of the hyperparameters permit to find subsets of features based on correlation measures. However, promoting diversity of reservoir measures does not necessarily correspond to the highest performance achievable. Thus, we adopted an evolutionary algorithm to better explore the space of possible combination of measurements (Algorithm 2).
It is necessary now to notice how a value of corresponds to a -dimensional boolean vector , whose -th dimension is zero if its -th feature is correlated more than with at least one other output. For each split in the outer loop, we adopted an evolutionary algorithm that operates over the -dimensional boolean space of feature-selection, where each individual corresponds to a specific vector . At each evolutionary step, we perform operations of crossover and mutation over a set of parents . For each split of the outer loop and at the first evolutionary step, we initialised the parents of the algorithm to . We defined a crossover operation among two individuals and as where the -th dimension of the new vector is randomly equal to or with the same probability. A mutation operation of a specific is defined as by simply applying the operator to each dimension of with a predefined probability . The application of crossovers and mutations permits the definition of a set of children from which we select the models with the highest performance over the test sets of the inner loop as parents for the next iteration. Otherwise stated, we selected the vectors corresponding to the lowest values of the average error as where selects the arguments of the corresponding function with minimal values. We notice how a step of the evolutionary approach aims to minimise an error estimated in the same fashion as in the algorithm of Algorithm 1, but this time searching for the best performing set , rather then the best performing couple of hyperparameter values and .
Finally, we stop the evolutionary algorithm at the iteration instance where the average performance of over is at minimum and selected the model with the lowest error on . The utilisation of a separate set for the stop-learning condition is necessary to avoid overfitting of the training data. Indeed, it is possible to notice how the performance on would improve for the first iterations of the evolutionary algorithm and then become worse. This concludes the evolutionary algorithm in Algorithm 2. At last, the overall performance of the model is computed as the sum of the mean-squared errors over the outer validation loop as . Summarising, we can think the overall methodology as an optimisation of relevant hyperparameters followed by a fine-tuning of the set of features used through an evolutionary algorithm. The final performance and its measure of variation reported in the paper are computed as average and standard deviation over ten repetitions of the evolutionary algorithm respectively.
Evaluating Overparameterisation
Meta-Learning
The goal of meta-learning is optimise an initial state for the network such that when a new task with limited data points is presented, the network can be quickly updated to give strong performance. A schematic of the meta-learning algorithm is presented in Figure 7 and pseudo-code shown in Algorithm 3.
Let us consider a family of M tasks . Each task is composed of a dataset and a cost function . A meta-learning algorithm is trained on a subset of and asked to quickly adapt and generalise on a new test subset of . Otherwise stated, the aim of meta-learning is to find an initial set of parameters that permits learning of an unseen task by updating the model over a small number of data points from . To achieve this, we use a variation of the MAML algorithm, which is now summarised. For a given task , the initial set of trainable parameters are updated via gradient descent over a batch of data points , where and are the inputs and targets for the j-th data point respectively
where the apex ′ is adopted to differentiate the data points used for the meta-update from the inner loop of Eq.1, and is the learning rate for the meta-update. Gradients of need to be computed with respect to , and this results in the optimisation of the recursive Eq.1 and the computation of higher-order derivatives. In our case, we use the first-order approximation of the algorithm[31]. After the meta-learning process, the system is asked to learn an unseen task updating through the iteration of Eq.1 computed on a small subset of data of .
Echo State Network Comparison
The Echo-state network of Figures 2 g-i) is a software model defined through
For the results of Figure 2, we evaluate the performance of a 100-node ESN. For each value of the dimensionality explored, we repeated the optimization process ten times resampling the random and . The black dots in Figure 2 report the average performance across these repetitions as the number of nodes varies. The black line reflects a polynomial fit of such results for illustrative purposes, while the grey area reflects the dispersion of the distributions of the results.
Multilayer Perceptron Comparison
Data availability statement
The experimental data used in this study is available at Github and Zenodo[32] under accession codes https://github.com/StenningK/NeuroOverParam.git and https://doi.org/10.5281/zenodo.12721639.
Code availability statement
The code developed in this study is available at Github and Zenodo[32] under accession codes https://github.com/StenningK/NeuroOverParam.git and https://doi.org/10.5281/zenodo.12721639.
Acknowledgements
This work was supported by the EPRSC grant EP/X015661/1 to WRB and HK
JCG was supported by the Royal Academy of Engineering under the Research Fellowship programme and the EPSRC ECR International Collaboration Grant ‘Three-Dimensional Multilayer Nanomagnetic Arrays for Neuromorphic Low-Energy Magnonic Processing’ EP/Y003276/1.
AV was supported by the EPSRC Centre for Doctoral Training in Advanced Characterisation of Materials (Grant No. EP/L015277/1) and EPSRC grant EP/X015661/1.
CC and TC performed the work as part of their Physics MSci project at Imperial College London.